Compute any interest accrued for each of the following notes payable owed by Penman, Inc., as of December 31, 2015 (use a 365-day year):


Lender

Issuance Date

Principal
Interest Rate (%)
Term
Nissim 11/21/2015 $20,000 12% 120 days
Klein 12/13/2015 16,000 9% 90 days
Bildersee 12/19/2015 18,000 10% 60 days

Round your answer to two decimal places.

Nissim Answer
Klein Answer
Bildersee Answer
Answer #1

Ans

Nissim $263.01
Klein $71.01
Bildersee $59.18

Explanation:

1) Nissim : 11/21 to 12/31 = 40 days

=$20000 × 12%= $2400

=$2400×40days/ 365 days

=$263.01

2) Klein: 12/13 to 12/31= 18 days

=$ 16000 ×9% =$1440

= $1440 × 18 days/ 365 days

= $ 71.01

3) Bildersee: 12/19 to 12/31= 12 days

= $ 18000 × 10% = $1800

=$1800 × 12days / 365 days

= $ 59.18